SOSU, Inc has before tax income this year is $900,000. The company’s payout ratio is 40%. The company's common equity currently has a book value of $5,000,000. They just paid a dividend of $1.87, and the required rate of return on this stock is 10%. Compute the value of this stock if dividends are expected to continue growing indefinitely at the company's internal growth rate. Tax rate = 28%.

 



1
Expert's answer
2021-04-15T07:58:07-0400

"Currentvalue=\frac{dividend in year 1}{Required rate-Growthrate}"

"Netincome=profit before tax\times (1-tax rate)"

"=900000\times(1-0.28)"

"=648,000"

"Growthrate=\frac{648,000}{5000000}\times100\times(1-0.40)=7.7767"

"Dividendin year1=1.87\times(1+0.0776)=2.01541"

"Value of the Stock=\frac{2.01541}{(10-7.776)}=\frac{2.01541}{2.224}=90.62"